Circular No. 7203

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1999 H1 (LEE) N. Biver, University of Hawaii; D. Bockelée-Morvan, P. Colom, J.

Crovisier, E. Gérard, and F. Henry, Observatoire de Paris-Meudon; J. K. Davies and H. E. Matthews, Joint Astronomy Centre; D. C. Lis and T. G. Phillips, California Institute of Technology (Caltech); A. Sievers and D. Teyssier, Institut de Radioastronomie Millimétrique (IRAM), Granada; L. Haikala, L. Nyman, and F. Rantakyro, Swedish-ESO Submillimetre Telescope (SEST), La Silla; and H. Weaver, Johns Hopkins University, report: “The HNC J(3-2) line at 272 GHz has been detected on June 6–7 at the James Clerk Maxwell Telescope (JCMT) with a line area of 0.07 ± 0.01 K km/s in brightness temperature. The HCN J(3-2) line observed on the same dates yields a HNC/HCN mixing ratio of 0.12, a value close to that found in comet C/1995 O1 at perihelion and twice higher than that inferred in C/1996 B2, which provides further constraints on the origin of HNC in comets. From coordinated observations undertaken at the SEST, Caltech Submillimeter Observatory, JCMT, IRAM 30-m, and Nançay radio telescopes from May 4 to June 13, several other species were observed, and their average abundances relative to water were derived as follows: HCN, 0.11 percent; CH_{3}OH, 4 percent; H_{2}CO, 1.3 percent; CS, 0.09 percent; CO, < 6 percent. The OH 18-cm transitions observed almost daily with the Nançay radio telescope provide water-production rates of 5.5 × 10^{28} s^{−1} during May 19–23 and 1.2 × 10^{29} s^{−1} during June 4–13.”

V382 VELORUM T. Napoleao, São Paulo, Brazil, reports that M. R. Mendes (Dois Cor-

regos, Brazil) obtained two prediscovery photographs of V382 Vel = N Vel 1999 (cf. IAUC 7184) with a 50-mm f/1.8 camera lens on Fuji Super G Plus (ISO 800) film as part of his monitoring of η Car. The nova was at the following estimated magnitudes on Mendes’ photographs: May 20.923 UT, ∼ 7.0–7.5 (20-s exposure); 21.944, ∼ 3.5 (30-s exposure).

Further selected visual magnitude estimates, some communicated via Napoleao: June 8.022 UT, 5.9 (A. S. Barros, Maceio, Brazil); 8.946, 6.3 (J. G. de S. Aguiar, Campinas, Brazil); 9.895, 6.3 (W. C. de Souza, São Paulo, Brazil); 10.456, 6.2 (A. Pearce, Nedlands, W. Australia); 11.15, 6.4 (V. Buso, Rosario, Argentina); 12.042, 6.7 (Napoleao); 13.436, 6.6 (Pearce); 14.948, 6.8 (J. Garcia, Rama Caida, Argentina); 15.448, 6.8 (Pearce); 15.917, 6.8 (Mendes).

1999 June 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green